Upon combustion, an unknown compound containing only carbon and hydrogen produces 55.50 g carbon dioxide and 22.72 g water. Determine the empirical formula of the unknown compound.

C2H3

CH3

C2H

CH2

CH
Answer #1

55.50 g carbon dioxide contains = 12 * 55.50 / 44 = 15.14 g carbon.

and

22.72 g water contains = 2 * 22.72 / 18 = 2.52 g

thus

mole ratio:

C : H

15.14 / 12 : 2.52 / 1.0

or

1.26 : 2.52

or

1 : 2

thus

empirical formula of the unknown compound is CH2.